OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own

OpenStudy (anonymous):

I just did this! haha

OpenStudy (anonymous):

lol its a pain in the butt right

OpenStudy (anonymous):

Yes! lol i can help you out though :)

OpenStudy (anonymous):

The first thing you're going to do is make a rational equation. Do you know how to do that?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

Otay, well the equation will be \[\frac{ 1 }{ 8 }+\frac{ 1 }{ 3 }=\frac{ 1 }{ x }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then, itll be \[\frac{ 8+3 }{ 24 }=\frac{ 1 }{ x }\]

OpenStudy (anonymous):

if t is the time it takes Jimmy to rake leaves alone, then t - 8 is the time it takes Harry to rake leaves alone. their rate together is 1/t + 1/(t-8). and given the both did in 3 hours 3[1/t + 1/(t-8)] = 1 t = 12 hours Answer is 12 hrs

OpenStudy (anonymous):

Add 8+3 and cross multiply and divide, itll be 12 like sourwing said

OpenStudy (anonymous):

omg thank you guys so much u help me out lot

OpenStudy (anonymous):

no problem :)

OpenStudy (radar):

This problem was closed, but I believe there were some misunderstandings regarding what the problem stated. One time for both working together was not 3 hours, Not the time for Jimmy to complete the job alone. Secondly, Harry can rake leaves 8 times faster than Jimmy, not 8 hours less than Jimmy.

OpenStudy (radar):

I may have some misunderstanding the problem myself. But I would have solved it this way. Let X = the time it takes Jimmy to do the job. Then X/8 = the time it takes Harry to do the job (Harry is 8 times faster than Jimmy) Together they can accomplish the job in 3 hours.